OpenStudy (anonymous):

Whats an equation i can use to graph coordinates. problem is y = x - 1

OpenStudy (anonymous):

make table of values

OpenStudy (anonymous):

find the y-intercept and then graph that along w/ some other values that you found using the slope

OpenStudy (anonymous):



OpenStudy (anonymous):

and how do i do that? i lost my notes so i dont know the equation to find them

OpenStudy (anonymous):

You should find x and y intercepts. x intercept is when y=0. So, x =1. y intercept is when x=0. So, y = -1. Now, draw line through two points (1,0) and (0,-1).

OpenStudy (anonymous):

how does x = 1 when there is no value to put in for x and y

OpenStudy (anonymous):

because of the way the equation is set up, the x and y intercepts will both have zero in them

OpenStudy (anonymous):

y = x -1 (when y = 0, you get 0 = x - 1, which means x = 1).

OpenStudy (anonymous):

y = x - 1 (when x = 0, you get y = 0 -1, which means y = -1).

OpenStudy (anonymous):

how do you know y is 0

OpenStudy (anonymous):

To draw a line, you just need two points. The two easiest points are the "intercepts". y is zero for the x intercept (where the line intersects x-axis). x is zero for the y intercept (where the line intersects y-axis).

OpenStudy (anonymous):

You don't know y is zero. You are forcing it to zero by focusing on the x-intercept.

OpenStudy (anonymous):